OpenStudy (anonymous):

What polynomial multiplyed by x-4 equals x^3 - 64??

OpenStudy (apoorvk):

x^3-64=x^3-4^3 now you know, a^3-b^3=(a-b)(a^2 +ab + b^2) so do you think you can compare and do it now??

OpenStudy (anonymous):

No, i am still confused

OpenStudy (anonymous):

\[\frac{x^3-64}{x-4}=x^2+4 x+16 \]